# Rigidity of anti-de Sitter (2+1)-spacetimes with convex boundary near the Fuchsian locus

### v1

@@ 2.4. Spacetimes with boundary and spaces of embeddings
Recall that we use GHC AdS (2+1)-spacetimes to mean globally hyperbolic compact anti-de Sitter (2+1)-spacetimes with spacelike convex boundary. Let $\mathcal{M}$ be the space of GHC AdS structures on $M$ with smooth strictly convex boundary. We pick up the viewpoint of locally homogeneous geometry, already described above, and consider elements of $\mathcal{M}$ as $G$-conjugacy classes of pairs $(\rho,f)$ of holonomies $\rho:\pi_{1}M=\pi_{1}S\rightarrow G$ and of $\rho$-equivariant developing maps $f:\widetilde{M}\rightarrow\mathbb{A}d\mathbb{S}^{3}$. We endow it with the topology of uniform $C^{\infty}$-convergence on compact subsets of $\widetilde{M}$. Let $\mathscr{M}$ be the quotient of $\mathcal{M}$ by isotopies. A structure is Fuchsian if its holonomy is conjugate into $G_{F}$.
We will, however, choose a different space to focus our consideration on. To this purpose, we will now notice that an element of $\mathscr{M}$ is determined by the structure of the boundary.
@@
We claim that $\mathscr{M}$ is naturally homeomorphic to $\mathscr{E}$. Indeed, from the description of the spaces it follows that there is a natural continuous injective map $\iota:\mathscr{M}\rightarrow\mathscr{E}$. It is also clear that $\iota$ is surjective. It only remains to check that $\iota^{-1}$ is continuous.
@@
We denote by $\mathcal{S}$ the space of Riemannian metrics on $S$ of curvature $<-1$, and denote by $\mathscr{S}$ its quotient by isotopies. We have the respective induced metric map $\mathcal{I}:\mathcal{E}\rightarrow\mathcal{S}\times\mathcal{S}$, descending to a map $\mathscr{E}\rightarrow\mathscr{S}\times\mathscr{S}$, which from now on we will denote by $\mathscr{I}$. We call Fuchsian elements of $\widehat{\mathcal{E}}$ whose representation is in $G_{F}$. This notion descends to elements of $\mathcal{E}$ and $\mathscr{E}$. We denote by ^$\widehat{\mathcal{E}}_{F}\subset\widehat{\mathcal{E}}$ the subspace of Fuchsian elements, and denote by $\mathcal{E}_{F}$ its quotient by conjugation. What we will actually prove is
@@
There exists a neighborhood $U$ of $\mathcal{E}_{F}$ in $\mathcal{E}$ such that the image of $U$ by $\mathcal{I}$ is open and $\mathcal{I}|_{U}$ is a homeomorphism onto its image.
@@ 2.5. Reminder on group cohomology
In the next section we discuss the differentiable structure on our spaces. We will then use basic properties of groups of representations, which we recall here. We briefly describe a construction that we will use several times. Let $\rho:\pi_{1}S\rightarrow G$ be a representation into a Lie group $G$ and $\mathfrak{m}$ be a $G$-module (most frequently we use $\mathfrak{m}=\mathfrak{g}$, though not only). It becomes a $\pi_{1}S$-module via the representation $\rho$. We denote by $Z^{1}_{\rho}(\pi_{1}S,\mathfrak{m})$ the vector space of (first) $\mathfrak{m}$-valued cocycles of $\pi_{1}S$, i.e., maps $\tau:\pi_{1}S\rightarrow\mathfrak{m}$ for all $\gamma_{1},\gamma_{2}\in\pi_{1}S$ satisfying
@@
Denote the subspace of coboundaries by $B^{1}_{\rho}(\pi_{1}S,\mathfrak{m})$ and define the (first) $\mathfrak{m}$-valued cohomology of $\pi_{1}S$ as
$H^{1}_{\rho}(\pi_{1}S,\mathfrak{m}):=Z^{1}_{\rho}(\pi_{1}S,\mathfrak{m})/B^{1}_{\rho}(\pi_{1}S,\mathfrak{m}).$
In particular, if $\widehat{\mathcal{R}}$ is the representation space of $\pi_{1}S$ into $G$, it is convenient to perceive its Zariski tangent space at $\rho$ as $Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$, where $\mathfrak{g}$ is considered as a $G$-module via the adjoint representation. This is a construction going back to Weil [63]. See also the book [52] for details. Then $B^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$ is the Zariski tangent space to the scheme associated to the $G$-orbit of $\rho$ under conjugation, and $H^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$ can be interpreted as the Zariski tangent space to the scheme associated to the quotient of $\widehat{\mathcal{R}}$ by conjugation of $G$. In our cases, the action of $G$ on the components of the representation space under our consideration is free and proper, and all these can be thought as the standard tangent spaces.
@@ 2.6. Differentiable structures on spaces of embeddings
As a main reference for Fréchet spaces and Fréchet manifolds we use an excellent survey [34] of Hamilton.
Consider a smooth path $(\rho_{t},f_{t})\in\widehat{\mathcal{E}}_{+}$ with $(\rho_{0},f_{0})=(\rho,f)$. As recalled in Section 2.5, a tangent vector to $\rho_{t}:\pi_{1}S\rightarrow G$ at $t=0$ can be described by a cocycle $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$, which is a map $\dot{\rho}:\pi_{1}S\rightarrow\mathfrak{g}$. A tangent vector to $f_{t}$ at $t=0$ is a vector field $\dot{f}$ defined on $f(\widetilde{S})$. An easy computation shows that the condition of $\rho_{t}$-equivariance of $f_{t}$ translates at the first order into the condition of automorphicity with respect to $\dot{\rho}$:
(2.3) $\dot{f}(\rho(\gamma)x)=d\rho(\gamma).(\dot{f}(x)+\dot{\rho}(\gamma)(x)).$
Here elements of $\mathfrak{g}$ are interpreted as Killing fields on $\mathbb{A}d\mathbb{S}^{3}$. An automorphic vector field is trivial if it is a restriction of a Killing field of $\mathbb{A}d\mathbb{S}^{3}$. Note that this in particular implies that $\dot{\rho}$ is a coboundary.
Pick $(\rho,f)\in\widehat{\mathcal{E}}_{+}$. Denote by $\widetilde{V}_{f}$ the restriction of $T\mathbb{A}d\mathbb{S}^{3}$ to $f(\widetilde{S})$. Denote by $C^{\infty}_{\rho}(\widetilde{V}_{f})$ the space of vector fields that are automorphic with respect to some $\dot{\rho}\in Z^{1}_{\rho}(\mathfrak{g})$. This is a Fréchet space, and $\widehat{\mathcal{E}}_{+}$ can be considered as a Fréchet manifold with charts in $C^{\infty}_{\rho}(\widetilde{V}_{f})$. In particular, $T_{(\rho,f)}\widehat{\mathcal{E}}_{+}\cong C^{\infty}_{\rho}(\widetilde{V}_{f})$. There is a natural fibration $\widehat{\pi}_{+}:\widehat{\mathcal{E}}_{+}\rightarrow\widehat{\mathcal{R}}$, i.e., a surjective submersion. The group $G$ acts on $\widehat{\mathcal{R}}$ freely and properly, and so it does on $\widehat{\mathcal{E}}_{+}$. The quotient $\mathcal{E}_{+}$ is also naturally a Fréchet manifold, and there is a surjective submersion $\pi_{+}:\mathcal{E}_{+}\rightarrow\mathcal{R}$. The same relates to the spaces $\widehat{\mathcal{E}}_{-}$, $\mathcal{E}_{-}$, $\widehat{\mathcal{E}}$ and $\mathcal{E}$. It is also standard that $\mathcal{S}$ is a Fréchet manifold.
@@ 2.7. Gauss-Codazzi fields on surfaces
Pick $s\in\mathcal{S}$. Thanks to the fundamental theorem of surfaces, the space of its isometric embeddings to $\mathbb{A}d\mathbb{S}^{3}$ has a nice parametrization in terms of GC-fields. Denote the curvature of $s$ by $\kappa:S\rightarrow\mathbb{R}$ and its connection by $\nabla$. A GC-field for $s$ is an $s$-self-adjoint operator field $b:TS\rightarrow TS$ satisfying the Gauss–Codazzi equations:
(1) $\det(b)=-\kappa-1$;
(2) $d^{\nabla}b=0$.
Denote by $\mathcal{GC}(s)\subset C^{\infty}(T^{1}_{1}S)$ the space of GC-fields on $(S,s)$. It is known from [59, Lemma 3.2] of Tamburelli that it is a finite-dimensional submanifold of dimension $66k-$ 6. Denote by $\mathcal{GC}_{+}(s)$ and $\mathcal{GC}_{-}(s)$ the subspaces of those $GC$-fields that have positive (resp. negative) eigenvalues. The fundamental theorem of surfaces (see, e.g., [3, Section 7]) allows to define two maps
$\psi_{h,\pm}:\mathcal{GC}_{\pm}(s)\rightarrow\mathcal{E}_{\pm}~{}$
so that a GC-field is the shape operator of the respective embedding. An examination of the proof of the fundamental theorem allows to ensure that $\psi_{h,\pm}$ are $C^{1}$-immersions. We will benefit from employing the dual viewpoint on the tangent vectors to $\mathcal{GC}_{\pm}(s)$: as tensor fields and as automorphic vector fields.
For a given $s\in\mathcal{S}$ with the almost-complex structure $j$ and a GC-field $b$ consider the following two metrics
(2.4) $s(({\rm id}+jb).,({\rm id}+jb).),~{}~{}~{}s(({\rm id}-jb).,({\rm id}-jb).)),$
@@
$b\mapsto\Big{(}\big{[}s(({\rm id}+jb).,({\rm id}+jb).)\big{]},\big{[}s(({\rm id}-jb).,({\rm id}-jb).))\big{]}\Big{)}.$
@@ 3. Infinitesimal rigidity
To this goal, in the next subsection we deal first with equivariant convex surfaces in Minkowski 3-space $\mathbb{R}^{2,1}$. We deduce from the work of Smith [58] that a pair of strictly future-convex and past-convex surfaces in $\mathbb{R}^{2,1}$, equivariant with respect to the same quasi-Fuchsian representation, is rigid with respect to the simultaneous automorphic infinitesimal isometric deformations. After that, in the next subsection we introduce the infinitesimal Pogorelov map from anti-de Sitter 3-space to Minkowski 3-space, and conclude Lemma 3.1 from the corresponding Minkowski result.
@@ 3.1. Equivariant surfaces in Minkowski 3-space
As previously, we call a representation $\rho:\pi_{1}S\rightarrow G^{0}$ quasi-Fuchsian if it is a deformation of a Fuchsian representation $\pi_{1}S\rightarrow G_{F}$. They form a component of a representation space. We denote by $\mathcal{R}^{0}$ the quotient of this component by conjugation of $G^{0}$, while we denote the component itself by ^$\widehat{\mathcal{R}}^{0}$. They and the associated spacetimes were actively studied also by Mess [44], as well as by Barbot, Bonsante, Seppi and others [4, 11, 6, 14].
For $\rho\in\widehat{\mathcal{R}}^{0}$ its linear part, i.e., the projection to the $G_{F}$-factor in (3.1), is a Fuchsian representation $\rho_{F}\in\widehat{\mathcal{R}}_{F}$. It is twisted by a cocycle $\tau\in Z^{1}_{\rho_{F}}(\pi_{1}S,\mathbb{R}^{2,1})$, where $\mathbb{R}^{2,1}$ is considered as a $G_{F}$-module via the standard action by isometries. This means that for $\gamma\in\pi_{1}S$ and $x\in\mathbb{R}^{2,1}$ we have
@@
The space ^$\widehat{\mathcal{R}}^{0}$ is a (trivial) vector bundle over ^$\widehat{\mathcal{R}}_{F}$ with the fiber at $\rho_{F}\in\widehat{\mathcal{R}}_{F}$ identified to $Z^{1}_{\rho_{F}}(\pi_{1}S,\mathbb{R}^{2,1})$. The conjugacy action of $G^{0}$, in particular, contains an action by coboundaries on each fiber, and $\mathcal{R}^{0}$ is a (trivial) vector bundle over $\mathcal{R}_{F}$ with the fiber at $\rho_{F}\in{\mathcal{R}}_{F}$ identified to $H^{1}_{\rho_{F}}(\pi_{1}S,\mathbb{R}^{2,1})$.
The Minkowski cross-product allows to construct a canonical isomorphism between $\mathbb{R}^{2,1}$ and $\mathfrak{g}_{F}$ as $G_{F}$-modules. See, e.g., [27, Section 3] of Fillastre–Seppi. Thereby, we have identification $H^{1}_{\rho_{F}}(\pi_{1}S,\mathfrak{g}_{F})\cong H^{1}_{\rho_{F}}(\pi_{1}S,\mathbb{R}^{2,1})$. Furthermore, we have
@@
We can repeat some definitions from Section 2.4 for immersions of surfaces into Minkowski 3-space. Let $s$ be a negatively curved metric on $S$ and $\kappa:S\rightarrow\mathbb{R}$ be its curvature. Denote by $\mathcal{GC}^{0}_{+}(s)$ and $\mathcal{GC}^{0}_{-}(s)$ the spaces of GC-fields on $(S,s)$ that have both eigenvalues of the respective sign. (For Minkowski 3-space the Gauss equation is $\det(b)=-\kappa$.) Similarly to the anti-de Sitter situation, they are endowed with topologies of smooth manifolds of dimension $66k-$ 6. The fundamental theorem of surfaces allows to interpret each element of $\mathcal{GC}_{\pm}(s)$ as an equivariant surface in $\mathbb{R}^{2,1}$. By considering the respective representation, we obtain two smooth maps
@@
If $b\in\mathcal{GC}^{0}_{\pm}(s)$, then the composition of $\phi_{s,\pm}(b)$ with the projection to $\mathscr{T}$ is given by the isotopy class of the metric $s(b.,b.)$, see, e.g., [58].
@@
The following result is a corollary of the work of Trapani–Valli [61], as it was noticed by Smith in [58].
@@
For any negatively curved $s$ the compositions of $\phi_{s,\pm}^{0}$ with the projection to $\mathscr{T}$ are diffeomorphisms.
In particular, the maps $\phi^{0}_{s,\pm}$ are immersions. Furthermore, it follows from Theorem 3.2 that every negatively curved metric $s$ determines via $\phi_{s,+}^{0}$ a vector field on $\mathscr{T}$, which we denote by $x_{s}$. If $f:\widetilde{S}\rightarrow\mathbb{R}^{2,1}$ is a $\rho$-equivariant future-convex isometric embedding for a quasi-Fuchsian $\rho$ with cocycle $\tau$, then $-f$ is a past-convex isometric embedding equivariant with respect to a quasi-Fuchsian representation with the same linear part and $-\tau$ as the cocycle. Hence, the same construction with the map $\phi_{s,-}^{0}$ produces the vector field $-x_{s}$.
There is another related vector field over $\mathscr{T}$ that we associate to $s$. Consider an equivariant isometric embedding $(\rho,f):\widetilde{S}\rightarrow\mathbb{R}^{2,1}$ with the induced metric $s$. Let $\rho_{F}$ be the linear part of $\rho$ and $\sigma$ be the hyperbolic metric on $S$ given by $\mathbb{H}^{2}/\rho_{F}$ and pulled to $S$ by the composition of the Gauss map from $f(\widetilde{S})$ and of $f$. Let $b$ be the shape operator of $f(\widetilde{S})$, pulled to $S$, and $a:=b^{-1}$. Note that $a$ is Codazzi for $\sigma$, see, e.g., [14, Proposition 2.7] of Bonsante–Seppi.
To proceed, we need to recall that due to the uniformization theorem, the Teichmüller space $\mathscr{T}$ can be also considered as the space of isotopy classes of conformal structures. See, e.g., [62] for details. A Riemannian metric $\sigma$ determines an isomorphism between the space of $\sigma$-symmetric positive-definite operator fields and the space of Riemannian metrics by $a\mapsto\sigma(a.,.)$. This allows to interpret the space of $\sigma$-symmetric operator fields as the tangent space at $\sigma$ to the space of Riemannian metrics. By taking the conformal class of a metric, we see that such an operator gives a tangent vector to $\mathscr{T}$ at the class of $\sigma$. We apply this to $\sigma$ and $a$ above. The result depends only on the isometry class of $(\rho,f)$, i.e., is determined by the shape operator $b\in\mathcal{GC}^{0}_{+}(s)$. Together with Theorem 3.2, this construction produces another vector field on $\mathscr{T}$, which we denote by $x^{\prime}_{s}$. Recall that there is a canonical complex structure on $\mathscr{T}$, which we denote by $\mathscr{J}$. Bonsante–Seppi showed that $\mathscr{J}x^{\prime}_{s}=x_{s}$ [14, Theorem B].

### v2

@@ 2.4. Spacetimes with boundary and spaces of embeddings
Recall that we use GHC AdS (2+1)-spacetimes to mean globally hyperbolic compact anti-de Sitter (2+1)-spacetimes with spacelike convex boundary. Let $\mathcal{M}$ be the space of GHC AdS structures on $M$ with smooth strictly convex boundary. We pick up the viewpoint of locally homogeneous geometry, already mentioned above, and consider elements of $\mathcal{M}$ as $G$-conjugacy classes of pairs $(\rho,f)$ of holonomies $\rho:\pi_{1}M=\pi_{1}S\rightarrow G$ and of $\rho$-equivariant developing maps $f:\widetilde{M}\rightarrow\mathbb{A}d\mathbb{S}^{3}$. We endow it with the topology of uniform $C^{\infty}$-convergence on compact subsets of $\widetilde{M}$. Let $\mathscr{M}$ be the quotient of $\mathcal{M}$ by isotopies. A structure is Fuchsian if its holonomy is conjugate into $G_{F}$.
We will, however, choose a different space to focus our consideration on. To this purpose, we will now notice that an element of $\mathscr{M}$ is determined by the “full” structure of the boundary, that is, by a lift of its two boundary components to equivariantly embedded complete surfaces in $\mathbb{A}d\mathbb{S}^{3}$. Our eventual goal in this paper is to show that, in certain cases, only a part of this data, namely the induced metrics, is sufficient to determine $\mathscr{M}$.
@@
We claim that $\mathscr{M}$ is naturally homeomorphic to $\mathscr{E}$. Indeed, from the description of the spaces it follows that there is a natural continuous injective map $\iota:\mathscr{M}\rightarrow\mathscr{E}$, sending a GHC AdS structure on $M$ with smooth and strictly convex boundary to the (equivalence class under $G$ of the) pair of equivariant embeddings $f_{\pm}:\widetilde{S}\to\mathbb{A}d\mathbb{S}^{3}$ of its boundary components. It is also clear that $\iota$ is surjective. It only remains to check that $\iota^{-1}$ is continuous.
@@
We denote by $\mathcal{S}$ the space of Riemannian metrics on $S$ of curvature <1<-1, and denote by $\mathscr{S}$ its quotient by isotopies. We have the respective induced metric map $\mathcal{I}:\mathcal{E}\rightarrow\mathcal{S}\times\mathcal{S}$, descending to a map $\mathscr{E}\rightarrow\mathscr{S}\times\mathscr{S}$, which from now on we will denote by $\mathscr{I}$. We call Fuchsian elements of $\widehat{\mathcal{E}}$ whose representation is in $G_{F}$. This notion descends to elements of $\mathcal{E}$ and $\mathscr{E}$. We denote by $\widehat{\mathcal{E}}_{F}\subset\widehat{\mathcal{E}}$ the subspace of Fuchsian elements, and denote by $\mathcal{E}_{F}$ its quotient by conjugation. What we will actually prove is
@@
There exists a neighborhood $U$ of $\mathcal{E}_{F}$ in $\mathcal{E}$ such that the image of $U$ by $\mathcal{I}$ is open, contains the diagonal and $\mathcal{I}|_{U}$ is a homeomorphism onto its image.
@@ 2.5. Reminder on group cohomology
In the next section, we discuss the differentiable structure on our spaces. We will then use basic properties of representation varieties, which we recall here. We briefly describe a construction that we will use several times. Let $\rho:\pi_{1}S\rightarrow G$ be a representation into a Lie group $G$ and $\mathfrak{m}$ be a $G$-module (most frequently we use $\mathfrak{m}=\mathfrak{g}$, though not only). It becomes a $\pi_{1}S$-module via the representation $\rho$. We denote by $Z^{1}_{\rho}(\mathfrak{m})=Z^{1}_{\rho}(\pi_{1}S,\mathfrak{m})$ the vector space of (first) $\mathfrak{m}$-valued cocycles of $\pi_{1}S$, i.e., maps $\tau:\pi_{1}S\rightarrow\mathfrak{m}$ for all $\gamma_{1},\gamma_{2}\in\pi_{1}S$ satisfying
@@
Denote the subspace of coboundaries by $B^{1}_{\rho}(\mathfrak{m})=B^{1}_{\rho}(\pi_{1}S,\mathfrak{m})$ and define the (first) $\mathfrak{m}$-valued cohomology of $\pi_{1}S$ as
$H^{1}_{\rho}(\mathfrak{m})=H^{1}(\pi_{1},\mathfrak{m}):=Z^{1}_{\rho}(\mathfrak{m})/B^{1}_{\rho}(\mathfrak{m}).$
In particular, if $\widehat{\mathcal{R}}$ is the representation variety of $\pi_{1}S$ into $G$, it is convenient to perceive its Zariski tangent space at $\rho$ as $Z^{1}_{\rho}(\mathfrak{g})$, where $\mathfrak{g}$ is considered as a $G$-module via the adjoint representation. This is a construction going back to Weil [67]. See also the book [55] for details. Then $B^{1}_{\rho}(\mathfrak{g})$ is the Zariski tangent space to the scheme associated to the $G$-orbit of $\rho$ under conjugation, and $H^{1}_{\rho}(\mathfrak{g})$ can be interpreted as the Zariski tangent space to the scheme associated to the quotient of $\widehat{\mathcal{R}}$ by conjugation of $G$. In our cases, the components of the representation variety under our consideration are smooth, the action of $G$ on it is free and proper, and all these can be thought as the standard tangent spaces.
@@ 2.6. Differentiable structures on spaces of embeddings
As a main reference for Fréchet spaces and Fréchet manifolds we use the excellent survey [36] of Hamilton.
Consider a smooth path $(\rho_{t},f_{t})\in\widehat{\mathcal{E}}_{+}$ with $(\rho_{0},f_{0})=(\rho,f)$. As recalled in Section 2.5, a tangent vector to $\rho_{t}:\pi_{1}S\rightarrow G$ at $t=0$ can be described by a cocycle $\dot{\rho}\in Z^{1}_{\rho}(\mathfrak{g})$, which is a map $\dot{\rho}:\pi_{1}S\rightarrow\mathfrak{g}$. A tangent vector to $f_{t}$ at $t=0$ is a vector field $\dot{f}$ defined on f(S~)f($\widetilde${S}). An easy computation shows that the condition of $\rho_{t}$-equivariance of $f_{t}$ translates at the first order into the condition of automorphicity with respect to $\dot{\rho}$:
(2.3) $\dot{f}(\rho(\gamma)x)=d\rho(\gamma)(\dot{f}(x))+\dot{\rho}(\gamma)(\rho(\gamma)x)~.$
Here elements of $\mathfrak{g}$ are interpreted as Killing fields on $\mathbb{A}d\mathbb{S}^{3}$. An automorphic vector field is trivial if it is the restriction of a Killing field of $\mathbb{A}d\mathbb{S}^{3}$. Note that this in particular implies that $\dot{\rho}$ is a coboundary.
Pick $(\rho,f)\in\widehat{\mathcal{E}}_{+}$. Denote by $\widetilde{V}_{f}$ the restriction of $T\mathbb{A}d\mathbb{S}^{3}$ to f(S~)f($\widetilde${S}). Denote by $C^{\infty}_{\rho}(\widetilde{V}_{f})$ the space of vector fields that are automorphic with respect to some $\dot{\rho}\in Z^{1}_{\rho}(\mathfrak{g})$. This is a Fréchet space, and $\widehat{\mathcal{E}}_{+}$ can be considered as a Fréchet manifold with charts in $C^{\infty}_{\rho}(\widetilde{V}_{f})$. In particular, $T_{(\rho,f)}\widehat{\mathcal{E}}_{+}\cong C^{\infty}_{\rho}(\widetilde{V}_{f})$. There is a natural fibration $\widehat{\pi}_{+}$$\widehat{\mathcal{E}}_{+}\rightarrow\widehat{\mathcal{R}}$, i.e., a surjective submersion. The group $G$ acts on $\widehat{\mathcal{R}}$ freely and properly, and so it does on $\widehat{\mathcal{E}}_{+}$. The quotient $\mathcal{E}_{+}$ is also naturally a Fréchet manifold, and there is a surjective submersion $\pi_{+}:\mathcal{E}_{+}\rightarrow\mathcal{R}$. The same relates to the spaces $\widehat{\mathcal{E}}_{-}$, $\mathcal{E}_{-}$, $\widehat{\mathcal{E}}$ and $\mathcal{E}$. It is also standard that $\mathcal{S}$ is a Fréchet manifold.
@@ 2.7. Gauss-Codazzi fields on surfaces
Pick $s\in\mathcal{S}$. Thanks to the fundamental theorem of surfaces, the space of its isometric embeddings to $\mathbb{A}d\mathbb{S}^{3}$ has a nice parametrization in terms of GC-fields. Denote the curvature of $s$ by $\kappa:S\rightarrow\mathbb{R}$ and its Levi-Civita connection by $\nabla$. A GC-field for $s$ is an $s$-symmetric operator field $b:TS\rightarrow TS$ satisfying the Gauss–Codazzi equations:
(Gauss) $\det(b)=-\kappa-1$;
(Codazzi) $d^{\nabla}b=0$.
Denote by $\mathcal{GC}(s)\subset C^{\infty}(T^{1}_{1}S)$ the space of GC-fields on $(S,s)$. It is known from [63, Lemma 3.2] of Tamburelli that it is a finite-dimensional submanifold of dimension 6k66k-6. Denote by $\mathcal{GC}_{+}(s)$ and $\mathcal{GC}_{-}(s)$ the subspaces of those $GC$-fields that have positive (resp. negative) eigenvalues. The fundamental theorem of surfaces (see, e.g., [3, Section 7]) allows to define two maps
$\psi_{s,\pm}:\mathcal{GC}_{\pm}(s)\rightarrow\mathcal{E}_{\pm}$~
so that a GC-field is the shape operator of the respective embedding. An examination of the proof of the fundamental theorem allows to ensure that
Theorem 2.2.
$\psi_{s,\pm}$ are $C^{1}$-immersions.
We will benefit from employing the dual viewpoint on the tangent vectors to $\mathcal{GC}_{\pm}(s)$: as tensor fields and as automorphic vector fields.
For a given $s\in\mathcal{S}$ with the almost-complex structure $j$ and a GC-field $b$, consider the following two metrics
(2.4) $s(({\rm id}+jb)\cdot,({\rm id}+jb)\cdot),~~~s(({\rm id}-jb)\cdot,({\rm id}-jb)\cdot)),$
@@
$b\mapsto\Big(\big[s(({\rm id}+jb)\cdot,({\rm id}+jb)\cdot)\big],\big[s(({\rm id}-jb)\cdot,({\rm id}-jb)\cdot))\big]\Big).$
@@ 3. Infinitesimal rigidity
To this goal, in the next subsection we deal first with equivariant convex surfaces in Minkowski 3-space $\mathbb{R}^{2,1}$. We deduce from the work of Smith [62] that a pair of strictly future-convex and past-convex surfaces in $\mathbb{R}^{2,1}$, equivariant with respect to the same quasi-Fuchsian representation, is rigid with respect to the simultaneous automorphic infinitesimal isometric deformations. After that, in the next subsection we introduce the infinitesimal Pogorelov map from anti-de Sitter 3-space to Minkowski 3-space, and conclude Proposition 3.1 from the corresponding Minkowski result.
@@ 3.1. Equivariant surfaces in Minkowski 3-space
As previously, we call a representation $\rho:\pi_{1}S\rightarrow G^{0}$ quasi-Fuchsian if it is a deformation of a Fuchsian representation $\pi_{1}S\rightarrow G_{F}$. They form a component of the representation variety. We denote by $\mathcal{R}^{0}$ the quotient of this component by conjugation of $G^{0}$, while we denote the component itself by $\widehat{\mathcal{R}}^{0}$. They and the associated spacetimes were actively studied also by Mess [46], as well as by Barbot, Bonsante, Seppi and others [4, 12, 6, 16].
For $\rho\in\widehat{\mathcal{R}}^{0}$, its linear part, i.e., the projection to the $G_{F}$-factor in (3.1), is a Fuchsian representation $\rho_{F}\in\widehat{\mathcal{R}}_{F}$. It is twisted by a cocycle $\tau\in Z^{1}_{\rho_{F}}(\mathbb{R}^{2,1})$, where $\mathbb{R}^{2,1}$ is considered as a $G_{F}$-module via the standard action by isometries. This means that for $\gamma\in\pi_{1}S$ and $x\in\mathbb{R}^{2,1}$ we have
@@
The space $\widehat{\mathcal{R}}^{0}$ is a (trivial) vector bundle over $\widehat{\mathcal{R}}_{F}$ with the fiber at $\rho_{F}\in\widehat{\mathcal{R}}_{F}$ identified to $Z^{1}_{\rho_{F}}(\mathbb{R}^{2,1})$. The conjugacy action of $G^{0}$, in particular, contains an action by coboundaries on each fiber, and $\mathcal{R}^{0}$ is a (trivial) vector bundle over $\mathcal{R}_{F}$ with the fiber at $\rho_{F}\in{\mathcal{R}}_{F}$ identified to $H^{1}_{\rho_{F}}(\mathbb{R}^{2,1})$.
The Minkowski cross-product allows to construct a canonical isomorphism between $\mathbb{R}^{2,1}$ and $\mathfrak{g}_{F}$ as $G_{F}$-modules. See, e.g., [28, Section 3] of Fillastre–Seppi. Thereby, we have identification $H^{1}_{\rho_{F}}(\mathfrak{g}_{F})\cong H^{1}_{\rho_{F}}(\mathbb{R}^{2,1})$. Furthermore, we have
@@
We can repeat some definitions from Section 2.4 for immersions of surfaces into Minkowski 3-space. Let $s$ be a negatively curved metric on $S$ and $\kappa:S\rightarrow\mathbb{R}$ be its curvature. Denote by $\mathcal{GC}^{0}_{+}(s)$ and $\mathcal{GC}^{0}_{-}(s)$ the spaces of GC-fields on $(S,s)$ that have both eigenvalues of the respective sign. (For Minkowski 3-space the Gauss equation is $\det(b)=-\kappa$.) Similarly to the anti-de Sitter situation, they are endowed with topologies of smooth manifolds of dimension 6k66k-6. The fundamental theorem of surfaces allows to interpret each element of $\mathcal{GC}^{0}_{\pm}(s)$ as a spacelike equivariant surface in $\mathbb{R}^{2,1}$. By considering the respective representation, we obtain two smooth maps
@@
Denote by $\chi_{s,\pm}^{0}:\mathcal{GC}^{0}_{\pm}(s)\rightarrow\mathscr{T}$ the compositions of $\phi_{s,\pm}^{0}$ with the projection to $\mathscr{T}$. For $b\in\mathcal{GC}^{0}_{\pm}(s)$, the element $\chi_{s,\pm}^{0}(b)$ is given by the isotopy class of the metric $s(b\cdot,b\cdot)$, see, e.g., [62].
@@
The following result is a corollary of the work of Trapani–Valli [65], as was noticed by Smith in [62].
@@
For any negatively curved metric $s$, $\chi_{s,\pm}^{0}$ are diffeomorphisms.
In particular, the maps $\phi^{0}_{s,\pm}$ are immersions. For $\sigma\in\mathscr{T}$ define
$x_{s}(\sigma):=\phi_{s,+}^{0}\circ(\chi_{s,+}^{0})^{-1}(\sigma).$
This gives us a vector field $x_{s}$ on $\mathscr{T}$ defined by $s$. If $f:\widetilde{S}\rightarrow\mathbb{R}^{2,1}$ is a $\rho$-equivariant future-convex isometric embedding for a quasi-Fuchsian $\rho$ with cocycle $\tau$, then f-f is a past-convex isometric embedding equivariant with respect to a quasi-Fuchsian representation with the same linear part and $\tau$ as the cocycle. Hence,
$x_{s}(\sigma)=\phi_{s,-}^{0}\circ(\chi_{s,-}^{0})^{-1}(\sigma).$
There is another related vector field over $\mathscr{T}$ that we associate to $s$. To proceed, we need to recall that due to the uniformization theorem, the Teichmüller space $\mathscr{T}$ can be also considered as the space of isotopy classes of conformal structures. See, e.g., [66] for details. Let $\mathcal{S}^{*}$ be the space of all Riemannian metrics on $S$. For $\sigma\in\mathcal{S}^{*}$ we denote by $\mathcal{O}_{\sigma}$ the space of $\sigma$-symmetric operator fields on $S$. Any $\sigma\in\mathcal{S}^{*}$ determines an isomorphism between the subspace of positive-definite elements of $\mathcal{O}_{\sigma}$ and $\mathcal{S}^{*}$ by $a\mapsto\sigma(a.,.)$. Denote the differential at the identity of this map by $\zeta_{\sigma}:\mathcal{O}_{\sigma}\rightarrow T_{\sigma}\mathcal{S}^{*}$, which is also an isomorphism. By taking the conformal class of a metric, we get the projection $\mathcal{S}^{*}\rightarrow\mathscr{T}$. We compose the differential of this projection at $\sigma$ with $\zeta_{\sigma}$ and obtain a map
$\eta_{\sigma}:\mathcal{O}_{\sigma}\rightarrow T_{\sigma}\mathscr{T}.$
Now define another vector field $x_{s}^{\prime}$ on $\mathscr{T}$ by
$x_{s}^{\prime}(\sigma):=\eta_{\sigma}\circ{\rm inv}\circ(\chi_{s,+}^{0})^{-1}(\sigma),$
where $\rm inv$} is the operation of taking the inverse on $C^{\infty}(T^{1}_{1}S)$, i.e., ${\rm inv}(b)=b^{-1}$.
Recall that there is a canonical complex structure on $\mathscr{T}$, which we denote by $\mathscr{J}$. Bonsante–Seppi showed that $\mathscr{J}x^{\prime}_{s}=x_{s}$ [16, Theorem B].